# Distributionally Robust Optimization with Adversarial Data Contamination

### v1

@@ 1 Introduction
Distributionally Robust Optimization (DRO) is a general stochastic optimization framework, where the goal is to minimize the loss of a target model on a worst-case distribution that is promised to lie in an ambiguity set—a set of distributions close to the reference distribution, with respect to a prespecified divergence. Intuitively, the ambiguity set models possible distributional shifts of the data. Since its introduction, the DRO framework has found a wide range of applications in a variety of contexts, including algorithmic fairness [hashimoto2018fairness] class imbalance [xu2020class], reinforcement learning [Lotidis2023, Yang2023, Wang2023a, Yu2023, Kallus2022, Liu2022, ramesh2024distributionally], robotics [Sharma2020], sparse neural network training [Sapkota2023], defense against model extraction [Wang2023b], and language model pretraining [Liu2021, xie2023doremi], finetuning [ma2025task], pruning [xia2023sheared, deng2024drpruning], and alignment [wu2025towards].
@@
Despite its advantages, the practical adoption of DRO is often hindered by the following vulnerability: its inherent sensitivity to outliers present in the training data; see, e.g., [hu2018does, zhu2022generalized]. This sensitivity is not merely a theoretical worst-case concern. Contaminated training data can significantly degrade the performance of DRO solutions, making it a critical barrier to its wider application in real-world scenarios where data quality is not always guaranteed [huang2023robust]. Notably, some DRO formulations, particularly those related to $f$-divergence-based measures (such as Conditional Value-at-Risk (CVaR)), provably assign higher importance to extreme data points [blanchet2024distributionally].
@@
The well-established field of robust statistics [Huber09, diakonikolas2022algorithmic] offers a natural and principled approach to modeling and mitigating the effects of data contamination in the training data. However, integrating these principles effectively within the DRO framework—in a manner that is both theoretically sound and computationally tractable—has turned out to be a considerable challenge. Many existing approaches that attempt to combine outlier robustness with distributional robustness struggle to maintain desirable properties or offer provable computational guarantees.
@@
A key desirable property of any outlier-robust DRO formulation is that it behaves sensibly under limiting conditions. In particular, it should satisfy two sanity checks: (1) when the ambiguity set, describing distributional shifts, reduces to a single distribution (i.e., the training and test distributions are identical and known, aside from potential contamination in the training data), the problem should naturally reduce to a standard outlier-robust supervised learning task; and (2) when there are no outliers in the training data, the formulation should recover the standard DRO problem for supervised learning. Many existing models fail to meet one or both of these criteria; see Section 3 for details.
@@
In this work, we initiate a systematic investigation of DRO in the presence of adversarial data contamination within the training set. As mentioned in the preceding discussion, in sharp contrast to prior formulations proposed with a similar motivation, our formulation captures both the DRO and the outlier-robust settings as special cases. We view this as an important conceptual contribution of this work. At the technical level, we aim to design the first statistically and computationally efficient algorithms for solving natural learning tasks robustly to distributional shifts, where an $\epsilon$-fraction of the training data can be adversarially corrupted according to the strong contamination model [diakonikolas2016robust-estimato].
@@
Given a parameter $20<{\epsilon}<1/$ 2 and an inlier distribution ${\mathbb{P}}_{0}$, an algorithm receives $\epsilon$-corrupted samples of size $N$ from ${\mathbb{P}}_{0}$ as follows: $N$ clean samples are first drawn i.i.d. from ${\mathbb{P}}_{0}$. An adversary is then allowed to inspect these samples, and replace an $\epsilon$-fraction of the samples with arbitrary points.
The strong contamination model is standard in robust statistics and subsumes other models like total variation contamination, where the contaminated samples are drawn i.i.d. from an unknown distribution ${{\mathbb{P}}_{\text{train}}}$ satisfying $\operatorname{TV}({\mathbb{P}}_{0},{{\mathbb{P}}_{\text{train}}})\leq{\epsilon}$.
@@
where we only have access to a dataset generated according to Definition 1.1. That is, we ask:
@@
Theorem 1.2 (Main Theorem–Informal; see Corollary 4.3).
Suppose the loss function takes the form of $\ell({\bm{w}};{\bm{x}},y)=\phi({\bm{w}}\cdot{\bm{x}},y)$ for some convex function $\phi$ that is $\zeta$-Lipschitz in the first coordinate for all $y$. There is a sample size $N=\tilde{O}(d/\epsilon)$ and an algorithm that, given an ${\epsilon}$-corrupted set of samples of size $N$, with high probability solves problem (1) approximately within $O(\|{\bm{w}}^{*}\|_{2}\sigma\zeta\sqrt{\epsilon})$ error, where ${\bm{w}}^{*}$ is the optimal solution.
We review related models for DRO with outliers in Section 3 and other related work in Appendix A.
@@ 1.1 Technical Overview
Modeling DRO with Adversarial Training Data Contamination
@@
We start by pointing out that effectively formulating DRO in the presence of adversarial training data contamination is subtle. Many prior approaches attempt to handle outliers in the training set by encoding information about potential contamination directly within the ambiguity set [nietert2023outlier], sometimes alongside modifications to the loss function [wang2024outlier]. While such models can, in principle, provide guarantees if the true data-generating distribution ${\mathbb{P}}_{0}$ remains within this augmented ambiguity set centered around the contaminated empirical distribution, they often suffer from various drawbacks. As discussed in Section 3.2 and further elaborated in Section C.1, these prior modeling strategies can lead to guarantees that depend undesirably on data dimensionality or may fail to recover standard robust estimation tasks in the limit of no distributional shift (i.e., when $\rho\to 0$). Moreover, as we argue in Section 3.1, such formulations often implicitly treat training data outliers in a way that is more aligned with handling outliers (or model misspecification) in the testing distribution.
@@
In sharp contrast, our modeling framework treats pre-decision (training data contamination) and post-decision (distributional shift at test time) uncertainties separately, aligned with [blanchet2024distributionally]. This separation allows for a more effective way to achieve robustness to both challenges simultaneously. We focus on the setting where an $\epsilon$-fraction of the training data is adversarially corrupted, and the goal is to find a model that performs well on a test distribution that may be shifted from the clean underlying data distribution ${\mathbb{P}}_{0}$.
Formulating and Solving DRO with Adversarial Training Data Contamination
@@
Having established our modeling principles, the challenge shifts to formulating and then efficiently solving the resulting optimization problem. The interplay between DRO and outlier-robustness presents several subtleties that require careful consideration. Employing $f$-divergence-based ambiguity sets, as explored in some prior work [zhai2021doro, bennouna2022holistic] combining DRO with outlier awareness, could be counterproductive in modeling DRO with training data contamination. Many $f$-divergences, and their corresponding DRO formulations (like those related to CVaR), are inherently sensitive to extreme data points, potentially amplifying the negative effect of outliers rather than mitigating it [blanchet2024distributionally].
@@
As our main technical contribution, we identify an expressive and tractable setting by focusing on Wasserstein-1 DRO ($W_{1}$-DRO) with convex Lipschitz loss functions of the form $\ell({\bm{w}};{\bm{x}},y)=\phi({\bm{w}}\cdot{\bm{x}},y)$, as outlined in Assumption 1.3. We note that the Lipschitz assumption of $\ell_{y_{i}}(\cdot)$ is a mild condition, because any loss function with superlinear growth admits +$\infty$+ ∞ worst-case loss under the setup of $W_{1}$-DRO, i.e., if $\ell_{y}(z)/|z|^{1+\iota}\to C$ as $|z|\to\infty$ for some $\iota>0$ and constant $C>0$, then $\sup_{{\mathbb{Q}}:W_{1}({\mathbb{P}},{\mathbb{Q}})\leq\rho}\mathbb{E}_{{\mathbb{Q}}}[\ell_{y}({\bm{w}}\cdot{\bm{x}})]=+\infty$ for all ${\bm{w}}$ and $\rho>0$ (see Fact B.3).
@@
This specific combination of a loss function and the ambiguity set is important for our algorithmic results. Prior work from the Wasserstein DRO literature [shafieezadeh2019regularization] allows us to reformulate the $W_{1}$-DRO problem (1) as an equivalent regularized risk minimization problem (3), which takes the empirical form (P)
@@
This step is crucial, as it converts the original min-max DRO problem into a minimization-only problem. However, the adversarial contamination in the training data means that we cannot directly compute empirical expectations or gradients using clean samples from ${\mathbb{P}}_{0}$. Instead, leveraging algorithmic results from robust statistics (Section 2), we show that the task effectively reduces to minimization of a generalized linear model with norm regularization, where access to the covariates is through an inexact oracle. This oracle processes the contaminated data and returns an estimate with bounded error $\Delta$, as formalized in Assumption 1.4 and demonstrated in Proposition 3.6.
@@
There exists $\Delta>0$ and an oracle that, on input an arbitrary $3\zeta$-uniformly bounded sequence ${\bm{\beta}}=(\beta^{i})_{i=1}^{N}$ and an ${\epsilon}$-corrupted version of a stable set (Fact 2.6) $\{{\bm{x}}^{s}\}_{s=1}^{N}$, outputs $\hat{{\bm{z}}}$ so that $\|\hat{\bm{z}}-\frac{1}{N}\sum_{i=1}^{N}\beta^{i}{\bm{x}}^{s}_{i}\|_{2}\leq\Delta$.
To efficiently solve this composite minimization problem with an inexact data oracle, we employ a primal-dual approach. By using Fenchel conjugacy (see Definition 2.2), we reformulate the objective to separate the non-linear but data-independent part of the loss function from the part that is linear in the data. This structural separation is vital, as it allows the robust mean estimation oracle to operate directly on appropriately weighted covariates to estimate the required linear term.
The core of our algorithmic approach is a variant of the Primal-Dual Hybrid Gradient (PDHG) method [chambolle2011first]. This algorithm is adapted to our setting with two main distinctions from standard PDHG:
@@
    The primal update step (Line 1 in Algorithm 1) relies on the inexact oracle from Proposition 3.6 to compute an estimate of the weighted sum of notionally clean covariates from the available contaminated samples.
@@
    The proximal term in the primal update includes a regularization parameter $c_{k}$ that strategically decreases over iterations, thus increasing primal steps as the algorithm progresses. This enables us to establish an estimation error guarantee (Proposition 4.1) that depends on the initial distance to the optimum, $\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}$, rather than on the potentially much larger diameter of the region containing all iterates.
A noteworthy feature of our algorithm and its analysis is that the robust mean estimation oracle is invoked only for the primal variable updates. In contrast, the dual variable updates (Line 1 in Algorithm 1) are performed using the raw, potentially contaminated, data points. Our convergence analysis carefully handles this asymmetry and the inexactness stemming from the oracle, by demonstrating that the algorithm’s iterates generated using contaminated data effectively track those of an idealized algorithm (Algorithm 2) that operates on underlying notionally clean (i.e., “stable”) data with a similar bounded-error oracle.
@@ 2 Preliminaries
Given a positive integer $N$, define $[N]:=\{1,2,\dots,N\}$. We use $\mathbb{I}_{{\mathcal{E}}}$ to denote the 0-1 indicator function of a set ${\mathcal{E}}$. We use $\chi$ to denote the characteristic function $\chi(0)=0$ and $\chi(a)=+\infty$ otherwise. For vectors ${\bm{x}}$ and $\widehat{\bm{x}}$ from the $d$-dimensional Euclidean space $\mathbb{R}^{d}$, we use $\langle{\bm{x}},\widehat{\bm{x}}\rangle$ and ${\bm{x}}\cdot\widehat{\bm{x}}$ to denote the standard inner product. We use $[x^{1},x^{2},\dots,x^{d}]^{\top}$ to denote the entries of ${\bm{x}}\in\mathbb{R}^{d}$. For $r>1$, the $\ell_{r}$ norm of a vector ${\bm{x}}\in\mathbb{R}^{d}$ is defined as $\|{\bm{x}}\|_{r}=(\sum_{i=1}^{d}|x^{i}|^{r})^{1/r}$. The $\ell_{\infty}$ norm is defined as $\|{\bm{x}}\|_{\infty}=\max_{1\leq i\leq d}|x^{i}|$. The operator norm of a matrix ${\bm{A}}\in\mathbb{R}^{m\times n}$, induced by the $\ell_{2}$ norm, is defined as $\|{\bm{A}}\|_{\operatorname{op}}=\sup_{\|{\bm{x}}\|_{2}=1}\|{\bm{A}}{\bm{x}}\|_{2}$. We write ${\bm{A}}\succeq{\bm{B}}$ to indicate that ${\bm{x}}^{\top}({\bm{A}}-{\bm{B}}){\bm{x}}\geq 0$ for all ${\bm{x}}\in\mathbb{R}^{d}$. For a set $S$ containing vectors of the same dimension, we let ${\bm{\mu}}_{S}$ denote their empirical mean and let ${\bm{\Sigma}}_{S}$ denote their empirical covariance matrix. For two functions $f$ and $g$, we say $f=\tilde{O}(g)$ if $f=O(g\log^{k}(g))$ for some constant $k$, and similarly define $\tilde{\Omega}$. We abuse the notation and use both $\sup$ and $\max$ to mean $\sup$.
Optimization

### v2

@@ 1 Introduction
Distributionally Robust Optimization (DRO) is a general stochastic optimization framework, where the goal is to minimize the loss of a target model on a worst-case distribution that is promised to lie in an ambiguity set—a set of distributions close to the reference distribution, with respect to a prespecified divergence. Intuitively, the ambiguity set models possible distributional shifts of the data. Since its introduction, the DRO framework has found a wide range of applications in a variety of contexts, including algorithmic fairness [Has+18a] class imbalance [Xu+20a], reinforcement learning [Lot+23a, Yan+23a, Wan+23b, Yu+23a, Kal+22a, Liu+22a, Sun+24a], robotics [Sha+20a], sparse neural network training [Sap+23a], defense against model extraction [Wan+23c], and language model pretraining [Liu+21a, Xie+23a], finetuning [Ma+25a], pruning [Xia+23a, Den+24a], and alignment [Wu+25a].
@@
Despite its advantages, the practical adoption of DRO is often hindered by the following vulnerability: its inherent sensitivity to outliers present in the training data; see, e.g., [Hu+18a, ZJS22a]. This sensitivity is not merely a theoretical worst-case concern. Contaminated training data can significantly degrade the performance of DRO solutions, making it a critical barrier to its wider application in real-world scenarios where data quality is not always guaranteed [Hua+23a]. Notably, some DRO formulations, particularly those related to $f$-divergence-based measures (such as Conditional Value-at-Risk (CVaR)), provably assign higher importance to extreme data points [Bla+25a].
@@
The well-established field of robust statistics [HR09a, DK23a] offers a natural and principled approach to modeling and mitigating the effects of data contamination in the training data. However, integrating these principles effectively within the DRO framework—in a manner that is both theoretically sound and computationally tractable—has turned out to be a considerable challenge. Many existing approaches that attempt to combine outlier robustness with distributional robustness struggle to maintain desirable properties or offer provable computational guarantees.
@@
A key desirable property of any outlier-robust DRO formulation is that it behaves sensibly under limiting conditions. In particular, it should satisfy two sanity checks: (1) when the ambiguity set, describing distributional shifts, reduces to a single distribution (i.e., the training and test distributions are identical and known, aside from potential contamination in the training data), the problem should naturally reduce to a standard outlier-robust supervised learning task; and (2) when there are no outliers in the training data, the formulation should recover the standard DRO problem for supervised learning. Many existing models fail to meet one or both of these criteria; see Section˜3 for details
@@
In this work, we initiate a systematic investigation of DRO in the presence of adversarial data contamination within the training set. As mentioned in the preceding discussion, in sharp contrast to prior formulations proposed with a similar motivation, our formulation captures both the DRO and the outlier-robust settings as special cases. We view this as an important conceptual contribution of this work. At the technical level, we aim to design the first statistically and computationally efficient algorithms for solving natural learning tasks robustly to distributional shifts, where an $\epsilon$-fraction of the training data can be adversarially corrupted according to the strong contamination model [Dia+16a].
@@
Given a parameter $0<{\epsilon}<1/2$ and an inlier distribution ${\mathbb{P}}_{0}$, an algorithm receives $\epsilon$-corrupted samples of size $N$ from ${\mathbb{P}}_{0}$ as follows: $N$ clean samples are first drawn i.i.d. from ${\mathbb{P}}_{0}$. An adversary is then allowed to inspect these samples and replace an $\epsilon$-fraction of the samples with arbitrary points.
The strong contamination model is standard in robust statistics and subsumes other models like total variation contamination, where the contaminated samples are drawn i.i.d. from an unknown distribution ${\mathbb{P}}_{\text{train}}$} satisfying [^1] The total variation distance between two distributions $\mathbb{P}$} and $\mathbb{Q}$} is defined as $\operatorname{TV}({\mathbb{P}},{\mathbb{Q}})=\sup_{A}|{\mathbb{P}}(A)-{\mathbb{Q}}(A)|$. $\operatorname{TV}({\mathbb{P}}_{0},{{\mathbb{P}}_{\text{train}}})\leq{\epsilon}$. Note here that, consistent with the literature in robust statistics, $\epsilon$} denotes the fraction of the outliers—not the target accuracy.
@@
where we only have access to a dataset generated according to Definition˜1.1. That is, we ask:
@@
Theorem 1.2 (Main Theorem–Informal; see Corollary˜4.5).
Suppose the loss function $\ell({\bm{w}};{\bm{x}},y)$ takes the form of $\phi({\bm{w}}\cdot{\bm{x}},y)$ for some convex function $\phi$ that is $\zeta$-Lipschitz in the first coordinate for all $y$. There is a sample size $N=\tilde{O}(d/\epsilon)$ and an algorithm that, given an $\epsilon$}-corrupted set of samples of size $N$, with high probability solves problem (1) approximately within $O(\|{\bm{w}}^{*}\|_{2}\sigma\zeta\sqrt{\epsilon})$ error, where $\bm{w}$$^{*}$ is the optimal solution and $\sigma$ is the variance of the covariates $\bm{x}$}.
The error guarantee in the above theorem scales naturally with the parameters of the problem, and we establish in Lemma˜4.8 that those dependencies are optimal. While the dependence on the initial distance to the optimum is necessary, we also show that it can be removed for specific problems like SVMs under stronger anti-concentration assumptions, as detailed in Lemma˜4.10.
@@ 1.1 Technical Overview
1.1.1 Modeling DRO with Adversarial Training Data Contamination
@@
We start by pointing out that effectively formulating DRO in the presence of adversarial training data contamination is subtle. Many prior approaches attempt to handle outliers in the training set by encoding information about potential contamination directly within the ambiguity set [NGS23a], sometimes alongside modifications to the loss function [Wan+24a]. While such models can, in principle, provide guarantees if the true data-generating distribution ${\mathbb{P}}_{0}$ remains within this augmented ambiguity set centered around the contaminated empirical distribution, they often suffer from various drawbacks. As discussed in Section˜3.2, these prior modeling strategies can lead to guarantees that depend undesirably on data dimensionality or may fail to recover standard robust estimation tasks in the limit of no distributional shift (i.e., when $\rho\to 0$). Moreover, as we argue in Section˜3.1, such formulations often implicitly treat training data outliers in a way that is more aligned with handling outliers (or model misspecification) in the testing distribution.
@@
In sharp contrast, our modeling framework treats pre-decision (training data contamination) and post-decision (distributional shift at test time) uncertainties separately, aligned with [Bla+25a]. This separation allows for a more effective way to achieve robustness to both challenges simultaneously. We focus on the setting where an $\epsilon$-fraction of the training data is adversarially corrupted, and the goal is to find a model that performs well on a test distribution that may be shifted from the clean underlying distribution ${\mathbb{P}}_{0}$.
1.1.2 Formulating and Solving DRO with Adversarial Training Data Contamination
@@
Having established our modeling principles, the challenge shifts to formulating and then efficiently solving the resulting optimization problem. The interplay between DRO and outlier-robustness presents several subtleties that require careful consideration. Employing $f$-divergence-based ambiguity sets, as explored in prior work [Zha+21a, BV22a] combining DRO with outlier awareness, could be counterproductive in modeling DRO with training data contamination. Many $f$-divergences, and their corresponding DRO formulations (like those related to CVaR), are inherently sensitive to extreme data points, potentially amplifying the negative effect of outliers rather than mitigating it [Bla+25a].
@@
As our main technical contribution, we identify an expressive and tractable setting by focusing on Wasserstein-1 DRO ($W_{1}$-DRO) with convex Lipschitz loss functions of the form $\ell({\bm{w}};{\bm{x}},y)=\phi({\bm{w}}\cdot{\bm{x}},y)$, as outlined in Assumption˜1.3. We note that the Lipschitz assumption of $\ell_{y_{i}}(\cdot)$ is a mild condition, because any loss function with superlinear growth admits $\infty$ worst-case loss under the setup of $W_{1}$-DRO, i.e., if $\ell_{y}(z)/|z|^{1+\iota}\to C$ as $|z|\to\infty$ for some $\iota>0$ and constant $C>0$, then $\sup_{{\mathbb{Q}}:W_{1}({\mathbb{P}},{\mathbb{Q}})\leq\rho}\mathbb{E}_{{\mathbb{Q}}}[\ell_{y}({\bm{w}}\cdot{\bm{x}})]=+\infty$ for all $\bm{w}$} and $\rho>0$ (see Fact˜3.1).
@@
This specific combination of a loss function and the ambiguity set is important for our algorithmic results. Prior work from the Wasserstein DRO literature [SKE19a] allows us to reformulate the $W_{1}$-DRO problem (1) as an equivalent regularized risk minimization problem (6), which takes the empirical form (P)
@@
This step is crucial, as it converts the original min-max DRO problem into a minimization-only problem. However, the adversarial contamination in the training data means that we cannot directly compute empirical expectations or gradients using clean samples from ${\mathbb{P}}_{0}$. Instead, leveraging algorithmic results from robust statistics (Section˜2), we show that the task effectively reduces to minimization of a generalized linear model with norm regularization, where access to the covariates is through an inexact oracle. This oracle processes the contaminated data and returns an estimate with bounded error $\Delta$, as formalized in Assumption˜1.4 and demonstrated in Proposition˜3.10.
@@
There exists $\Delta>0$ and an oracle that, on input an arbitrary $3\zeta$-uniformly bounded sequence ${\bm{\beta}}=(\beta^{i})_{i=1}^{N}$ and an $\epsilon$}-corrupted version of a stable set (Fact˜2.6) $\{{\bm{x}}^{s}\}_{s=1}^{N}$, outputs $\hat{{\bm{z}}}$ so that $\|\hat{\bm{z}}-\frac{1}{N}\sum_{i=1}^{N}\beta^{i}{\bm{x}}^{s}_{i}\|_{2}\leq\Delta$.
To efficiently solve this composite minimization problem with an inexact data oracle, we employ a primal-dual approach. By using Fenchel conjugacy (see Definition˜2.2), we reformulate the objective to separate the non-linear but data-independent part of the loss function from the part that is linear in the data. This structural separation is vital, as it allows the robust mean estimation oracle to operate directly on appropriately weighted covariates to estimate the required linear term.
The core of our algorithmic approach is a variant of the Primal-Dual Hybrid Gradient (PDHG) method [CP11a]. This algorithm is adapted to our setting with two main distinctions from standard PDHG:
@@
    The primal update step (Line 1 in Algorithm˜1) relies on the inexact oracle from Proposition˜3.10 to compute an estimate of the weighted sum of notionally clean covariates from the available contaminated samples.
@@
    The proximal term in the primal update includes a regularization parameter $c_{k}$ that strategically decreases over iterations, thus increasing primal steps as the algorithm progresses. This enables us to establish an estimation error guarantee (Proposition˜4.1) that depends on the initial distance to the optimum, $\|{\bm{w}}_{0}-{\bm{w}}^{*}\|_{2}$, rather than on the potentially much larger diameter of the region containing all iterates.
A noteworthy feature of our algorithm and its analysis is that the robust mean estimation oracle is invoked only for the primal variable updates. In contrast, the dual variable updates (Line 1 in Algorithm˜1) are performed using the raw, potentially contaminated, data points. Our convergence analysis carefully handles this asymmetry and the inexactness stemming from the oracle, by demonstrating that the algorithm’s iterates generated using contaminated data effectively track those of an idealized algorithm (Algorithm˜2) that operates on underlying notionally clean (i.e., “stable”) data with a similar bounded-error oracle.
@@ 2 Preliminaries
Given a positive integer $N$, define $N]:=\{1,2,\dots,N\}$. We use $\mathbb{I}_{{\mathcal{E}}}$ to denote the 0-1 indicator function of a set $\mathcal{E}$}. We use $\chi$ to denote the characteristic function $\chi(0)=0$ and $\chi(a)=+\infty$ otherwise. For vectors $\bm{x}$} and $\widehat{\bm{x}}$ from the $d$-dimensional Euclidean space $\mathbb{R}^{d}$, we use $\langle{\bm{x}},\widehat{\bm{x}}\rangle$ and $^{\bm{x}}\cdot\widehat{\bm{x}}$ to denote the standard inner product. We use $x^{1},x^{2},\dots,x^{d}]^{\top}$ to denote the entries of ${\bm{x}}\in\mathbb{R}^{d}$. For $r>1$, the $\ell_{r}$ norm of a vector ${\bm{x}}\in\mathbb{R}^{d}$ is defined as $\|{\bm{x}}\|_{r}=(\sum_{i=1}^{d}|x^{i}|^{r})^{1/r}$. The $\ell_{\infty}$ norm is defined as $\|{\bm{x}}\|_{\infty}=\max_{1\leq i\leq d}|x^{i}|$. The operator norm of a matrix ${\bm{A}}\in\mathbb{R}^{m\times n}$, induced by the $\ell_{2}$ norm, is defined as $\|{\bm{A}}\|_{\operatorname{op}}=\sup_{\|{\bm{x}}\|_{2}=1}\|{\bm{A}}{\bm{x}}\|_{2}$. We write ${\bm{A}}\succeq{\bm{B}}$ to indicate that ${\bm{x}}^{\top}({\bm{A}}-{\bm{B}}){\bm{x}}\geq 0$ for all ${\bm{x}}\in\mathbb{R}^{d}$. For a set $S$ containing vectors of the same dimension, we let ${\bm{\mu}}_{S}$ denote their empirical mean and let ${\bm{\Sigma}}_{S}$ denote their empirical covariance matrix. For two functions $f$ and $g$, we say $f=\tilde{O}(g)$ if $f=O(g\log^{k}(g))$ for some constant $k$, and similarly define $\tilde{\Omega}$.
2.1 Optimization